# SSF-Net: Spatial-Spectral Fusion Network with Spectral Angle Awareness for Hyperspectral Object Tracking

### v1

@@ V Conclusion
In this article, we proposed a spatial-spectral fusion network with spectral angle awareness (SSF-Net) for hyperspectral object tracking. We first designed spatial-spectral feature backbone $S^{2}$ to extract the more robust spatial and spectral features compared with the state-of-the-art trackers. Then, a spectral attention fusion module (SAFM) was proposed to incorporate the visual information from the RGB modality into the HS modality to obtain robust HS fusion features. Finally, a spectral angle awareness module (SAAM) was designed to predict the more accurate position of objects. With the proposed spectral angle awareness loss (SAAL), the tracker can be trained in an end-to-end way. Experimental results on the HOTC dataset showed the effectiveness of our proposed methods.

### v2

@@ V Conclusion
In this article, we proposed a spatial-spectral fusion network with spectral angle awareness (SSF-Net) for hyperspectral object tracking. We first designed a spatial-spectral feature backbone $S^{2}$ to extract more robust spatial and spectral features. Then, a spectral attention fusion module (SAFM) was proposed to incorporate the visual information from the RGB modality into the HS modality to obtain robust HS fusion features. Finally, a spectral angle awareness module (SAAM) was designed to predict the more accurate position of objects. With the proposed spectral angle awareness loss (SAAL), the tracker can be trained in an end-to-end way. Experimental results on the HOTC-2020, HOTC-2024, and BihoT dataset showed the effectiveness of our proposed methods and its superiority over the state-of-the-arts.